OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 0) and (2, 8)? y = 2(x + 2) y = 2(x – 2) y = –2(x – 2) y = –2(x + 2)

OpenStudy (anonymous):

point-slope form: y1 - y2 = m (x1 - x2) Then, find the slope of the line. slope equation: y1 - y2 / x1 - x2 8 - 3 / 2 - 3 5 / -1 -5 Next, plug in the coordinates in the right place. x2: 2 or 3 y2: 8 or 4 plug-in: y - 4 = -5 (x - 3) y - 8 = -5 (x - 2) And that's the equation you can use either coordinates if you'd like.

OpenStudy (anonymous):

sorry I am doing my HW and giving you your answers

OpenStudy (anonymous):

so needed to take it frm there

OpenStudy (anonymous):

y = 2(x – 2)

OpenStudy (mimi_x3):

\(y=mx+b\) Use the gradient formula to find \(m\) which is the gradient: \[m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}} \] Then sub then \(x-value\) and \(y-value\) to find \(b\) Then you will find the equation. Try it!